# LLM-based relevance assessment still can’t replace human relevance assessment

### v1

@@ 5 When Automatic Judgments become Useless
While the uwc1 run demonstrates how a bad actor can strategically subvert an evaluation experiment, it is reasonable to assume that most participants are well-intentioned. These participants are not merely competing to win but are contributing to the creation of reusable test collections that support the development of innovative systems. However, even without malicious intent, the next generation of information systems will likely incorporate the latest advancements in LLMs, including prompting for relevance. As a result, some ranking methods will inherently embed elements that mirror LLM-based relevance judgments. In a future evaluation experiment, it is plausible that even a well-intentioned participant could inadvertently undermine the evaluation process.
@@
Looking ahead, we anticipate that future retrieval systems will increasingly rely on automatically generated training data to optimize machine learning components. Here, Goodhart’s law serves as a cautionary principle: “When a measure becomes a target, it ceases to be a good measure.” While the current observed correlation between manual and automatic assessment methods is strong, we predict that this correlation will degrade as developers incorporate LLM-based evaluation components into their systems. Over time, these systems risk becoming disconnected from the human judgments they are intended to serve. If the entire end-to-end experimental pipeline – from query formulation to relevance labeling – is fully automated, the evaluation process devolves into an LLM assessing its own assessments. The circularity feared by Faggioli et al. [5] is no longer a hypothetical concern; it has already begun to manifest in practice.
6 Conclusion
@@
This paper raises serious concerns about the claims made by Upadhyay et al. [8], which presents a preliminary analysis of data from the retrieval (“R”) task of the TREC 2024 RAG Track. The author list of Upadhyay et al. [8] includes some of the most prominent experts in the area of information retrieval evaluation. Despite being preliminary, their conclusions strongly imply that LLM-based relevance assessment can replace human relevance assessment — a claim that does withstand scrutiny. Given the authority of the authors and the strength of their implied conclusions, there is a risk that these findings may gain widespread acceptance within the research community without sufficient critical consideration. Nearly two years ago, Faggioli et al. [5] reached the opposite conclusion based on similar evidence. Their concerns have still not been addressed.
@@ Acknowledgments
We have discussed our concerns with some of the authors of Upadhyay et al. [8], and we appreciate their attention and feedback. They also generously provided us with access to their non-public data to confirm factual statements in this paper.

### v2

@@ 6. When Automatic Judgments become Useless
While the uwc1 run demonstrates how a bad actor can strategically subvert an evaluation experiment, it is reasonable to assume that most participants are well-intentioned. These participants are not merely competing to win but are contributing to the creation of reusable test collections that support the development of innovative systems. However, even without malicious intent, the next generation of information systems will likely incorporate the latest advancements in LLMs, including prompting LLMs for relevance. As a result, some ranking methods will inherently embed elements that mirror LLM-based relevance judgments. In a future evaluation experiment, it is plausible that even a well-intentioned participant could inadvertently undermine the evaluation process.
@@
Looking ahead, we anticipate that future retrieval systems will increasingly rely on automatically generated training data to optimize machine learning components. Here, Goodhart’s law serves as a cautionary principle (Goodhart, 1975): “When a measure becomes a target, it ceases to be a good measure.” While the current observed correlation between manual and automatic assessment methods is strong, we predict that this correlation will degrade as developers incorporate LLM-based evaluation components into their systems in more refined ways than our simulation in Section 4.3. Over time, these systems risk becoming disconnected from the human judgments they are intended to serve. If the entire end-to-end experimental pipeline — from query formulation to relevance labeling — is fully automated, the evaluation process devolves into an LLM assessing its own assessments. The circularity feared by Faggioli et al. (2023) is no longer a hypothetical concern; it has already begun to manifest in practice.
7. Conclusion
@@
This paper raises serious concerns about the claims made by Upadhyay et al. (2024a), which presents a preliminary analysis of data from the retrieval (“R”) task of the TREC 2024 RAG Track. The author list of Upadhyay et al. (2024a) includes some of the most prominent experts in the area of information retrieval evaluation. Despite being preliminary, their conclusions strongly imply that LLM-based relevance assessment can replace human relevance assessment — a claim that does not withstand scrutiny. Given the authority of the authors and the strength of their implied conclusions, there is a risk that these findings may gain widespread acceptance within the research community without sufficient critical consideration. Nearly two years ago, Faggioli et al. (2023) reached the opposite conclusion based on similar evidence. Their concerns have still not been addressed.
@@ Acknowledgments
We have discussed our concerns with some of the authors of Upadhyay et al. (2024a), and we appreciate their attention and feedback. They also generously provided us with early access to their data to confirm factual statements in this paper.
This material is based in part upon work supported by the National Science Foundation under Grant No. 1846017. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the National Science Foundation.